# ActiveGS: Active Scene Reconstruction using Gaussian Splatting

### v1

@@ Abstract
Robotics applications often rely on scene reconstructions to enable downstream tasks. In this work, we tackle the challenge of actively building an accurate map of an unknown scene using an on-board RGB-D camera. We propose a hybrid map representation that combines a Gaussian splatting map with a coarse voxel map, leveraging the strengths of both representations: the high-fidelity scene reconstruction capabilities of Gaussian splatting and the spatial modelling strengths of the voxel map. The core of our framework is an effective confidence modelling technique for the Gaussian splatting map to identify under-reconstructed areas, while utilising spatial information from the voxel map to target unexplored areas and assist in collision-free path planning. By actively collecting scene information in under-reconstructed and unexplored areas for map updates, our approach achieves superior Gaussian splatting reconstruction results compared to state-of-the-art approaches. Additionally, we demonstrate the applicability of our active scene reconstruction framework in the real world using an unmanned aerial vehicle.
@@
L. Jin, X. Zhong, Y. Pan, J. Behley, C. Stachniss are with the Center for Robotics, University of Bonn, Germany. M. Popović is with the MAVLab, TU Delft, the Netherlands. C. Stachniss is also with the Lamarr Institute for Machine Learning and Artificial Intelligence, Germany. This work has been fully funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy, EXC-2070 – 390732324 (PhenoRob). Corresponding author: ljin@uni-bonn.de.
@@ I Introduction
In this work, we tackle the problem of actively reconstructing unknown scenes using posed RGB-D camera data. Given a limited budget, e.g. mission time, our goal is to obtain an accurate 3D representation of the scene by actively positioning the robot with its camera online during a mission. Existing active scene reconstruction frameworks mainly rely on conventional map representations such as voxel grids, meshes, or point clouds [16, 31, 30, 2, 29, 39, 38]. However, these methods often do not deliver high-fidelity reconstruction results due to their sparse representations. Recent advances in implicit neural representations, e.g. neural radiance fields (NeRFs) [20], have attracted significant research interest for their accurate dense scene reconstruction capabilities and low memory footprints. In the context of active reconstruction, several emerging works [24, 5, 13, 26, 36] incorporate uncertainty estimation in NeRFs and exploit it to guide view planning. While these approaches demonstrate promising results, the rather costly volumetric rendering procedure during online incremental mapping poses limitations for NeRF-based active scene reconstruction.
@@
We make the following three claims: (i) our ActiveGS framework achieves superior reconstruction performance compared to state-of-the-art NeRF-based approach and GS-based baselines; (ii) our explicit confidence modelling for Gaussian primitives enables informative viewpoint evaluation and targeted inspection around under-reconstructed surfaces, further improving mission efficiency and reconstruction quality; (iii) we validate our approach in different synthetic indoor scenes and in a real-world scenario using an unmanned aerial vehicle (UAV). To support reproducibility and future research, our implementation will be available at: https://github.com/dmar-bonn/active-gs.
@@ II-B Active Scene Reconstruction
To address this, recent research explores implicit neural representations for active reconstruction. In an object-centric setup, several methods incorporate uncertainty estimation into NeRFs [24, 13, 36, 26] and use this information to select next best viewpoints. For scene-level reconstruction, Yan et al. [37] and Kuang et al. [17] investigate the loss landscape of implicit neural representations during training to identify under-reconstructed areas. NARUTO [5] learns an uncertainty grid map alongside a hybrid neural scene representation, guiding data acquisition in uncertain regions. These implicit neural representations often face challenges such as inefficient map updates and catastrophic forgetting during incremental mapping.
Several concurrent works propose using GS for active scene reconstruction. GS-Planner [14] and HGS-planner [34] incorporate unknown voxels into the GS rendering pipeline and detect unseen regions for exploration. Li et al. [18] uses a Voronoi graph to extract a traversable topological map from the GS map for efficient path planning. However, this approach is limited to a 2D planning space, reducing its effectiveness. FisherRF [12] evaluates the information content of a novel view by measuring the Fisher information value in the GS parameters. This procedure requires computationally expensive gradient calculations at each training and candidate viewpoint, making it inefficient for active scene reconstruction. We build upon the idea of using GS for active scene reconstruction, while introducing a key innovation: we explicitly model the confidence of each Gaussian primitive, enabling viewpoint sampling around low-confidence Gaussian primitives for targeted inspection and fast feed-forward confidence rendering for informative viewpoint evaluation.
@@ III Our Approach
Figure 2: Overview of our ActiveGS framework. We utilise a 2D GS map for high-fidelity scene reconstruction and a coarse voxel map for exploration and path planning. Given posed RGB-D measurements as input, we incrementally update our map representation. Our adaptive view planner leverages unexplored region information in the voxel map for exploration while exploiting low-confidence Gaussian primitives to enable targeted inspection. We iterate between the map update and view planning steps until a pre-allocated mission time is reached.
@@ III-E Viewpoint Sampling and Evaluation
Our viewpoint sampling strategy involves two types of candidate viewpoints. First, we randomly sample $N_{\text{random}}$ candidate viewpoints within a specified range around the current viewpoint. However, relying solely on random local sampling can lead to local minima. To address this, we introduce additional candidate viewpoints based on regions of interest (ROI) defined in the voxel map. We begin by identifying frontier voxels [35] and add them to our ROI set $\mathcal{R}$. By explicitly modelling the confidence of each Gaussian primitive, we can identify and also include voxels containing low-confidence Gaussian primitives in $\mathcal{R}$. For each voxel in $\mathcal{R}$, we define normals indicating the best viewing direction. For voxels with low-confidence Gaussian primitives, this is simply the average normal of these Gaussian primitives. The normal of frontier voxels is determined by finding their neighbouring free voxels and calculating the average directional vector from the frontier voxel to these neighbours. To generate ROI-based candidate viewpoints, we create a fixed number of candidate viewpoints within the cone defined by the minimum and maximum sampling distances to each ROI’s centre, and the maximum angular difference relative to its normal. Starting from the closest ROI voxel, we continue outward until we have collected up to $N_{\text{ROI}}$ viewpoints in free space. We illustrate the sampling process in Fig. 3.
@@ IV-A Implementation Details
We test our implementation on a desktop PC with an Intel Core i9-10940X CPU and an NVIDIA RTX A5000 GPU. In this setup, one mapping and planning steps take approximately $1$ s and $0.5$ s, respectively. The whole framework consumes $4-5$ GB GPU RAM during an online mission, depending on the size of the scene.
@@ IV-B Simulation Experiments
  • Ours (w/o ROI): A variant of our ActiveGS that considers only local random sampling, with $N_{\text{ROI}}=0$.
  • FBE [35]: Frontier-based exploration framework for covering unexplored regions in the scene. We use the collected RGB-D data for updating GS map, similar to our framework.
@@
We present the results of simulation experiments in Fig. 4. Our approach achieves the best performance in both rendering and mesh quality in all test scenes, supporting our first claim that our approach outperforms state-of-the-art NeRF and GS-based methods. The NeRF-based active scene reconstruction framework, NARUTO, exhibits a significant performance gap compared to our approach, particularly in RGB rendering. This disparity arises because NeRF-based approaches often compromise model capacity to enable faster map updates, bottlenecking their representation quality in scene-level reconstruction. FisherRF evaluates the utility of a viewpoint by calculating the Fisher information contained in the parameters of the Gaussian primitives within its field of view. This requires computationally expensive gradient calculation for all candidate and training viewpoints, leading to prolonged planning times and incomplete reconstruction under limited mission time. Moreover, since Fisher information is conditioned on the candidate viewpoint, the viewpoint must be selected before its utility can be evaluated. Consequently, viewpoint sampling cannot be directly informed by Fisher information. In contrast, our approach models the confidence of Gaussian primitives, enabling fast feed-forward confidence rendering for viewpoint evaluation and explicit identification of surfaces with low confidence. This allows us to generate candidate viewpoints targeted at under-reconstructed areas, significantly enhancing the reconstruction quality and efficiency. FBE, on the other hand, does not consider surface reconstruction quality during online missions, limiting the potential for further quality improvement after exploration. The ablation study comparing Ours and Ours (w/o ROI) demonstrates that our approach benefits from ROI-based sampling for targeted inspection, as reflected by the higher mean and smaller standard deviation in both evaluation metrics. These results confirm that our confidence modelling is effective in achieving efficient and high-fidelity active scene reconstruction, validating our second claim. We visually compare the reconstruction results in Fig. 5.
@@ IV-C Real-World Experiments
We validate our framework in a real-world experiment using a UAV equipped with an Intel RealSense 455 RGB-D camera to reconstruct a scene of size $6$ m $\times$ $6$ m $\times$ $3$ m. Unlike simulation experiments, we do not account for the pitch angle of viewpoints in this experiment due to control limitations. The UAV pose is tracked by an OptiTrack motion capture system. Given the limited on-board resources, we run ActiveGS on our desktop PC, where it receives RGB-D and pose data from the UAV for map updates and sends planned collision-free waypoints to guide the UAV. All communication is handled via ROS [25].
@@ V Conclusions and Future Work
One limitation of our current framework is that it considers perfect localisation. To address this, future work will incorporate the localisation uncertainty into the confidence modelling of Gaussian primitives. Additionally, we will explore better ways of integrating the voxel map and GS map to achieve more efficient online mapping and view planning.

### v2

@@ Abstract
Robotics applications often rely on scene reconstructions to enable downstream tasks. In this work, we tackle the challenge of actively building an accurate map of an unknown scene using an RGB-D camera on a mobile platform. We propose a hybrid map representation that combines a Gaussian splatting map with a coarse voxel map, leveraging the strengths of both representations: the high-fidelity scene reconstruction capabilities of Gaussian splatting and the spatial modelling strengths of the voxel map. At the core of our framework is an effective confidence modelling technique for the Gaussian splatting map to identify under-reconstructed areas, while utilising spatial information from the voxel map to target unexplored areas and assist in collision-free path planning. By actively collecting scene information in under-reconstructed and unexplored areas for map updates, our approach achieves superior Gaussian splatting reconstruction results compared to state-of-the-art approaches. Additionally, we demonstrate the real-world applicability of our framework using an unmanned aerial vehicle.
Index Terms:
Mapping; RGB-D Perception
@@
Manuscript received: December 20, 2024; Revised: February 23, 2025; Accepted: March 20, 2025.This paper was recommended for publication by Editor Javier Civera upon evaluation of the Associate Editor and Reviewers’ comments. This work has been fully funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy, EXC-2070 – 390732324 (PhenoRob).1L. Jin, X. Zhong, Y. Pan, J. Behley, C. Stachniss are with the Center for Robotics, University of Bonn, Germany. 2M. Popović is with the MAVLab, TU Delft, the Netherlands. 3C. Stachniss is also with the Lamarr Institute for Machine Learning and Artificial Intelligence, Germany.Digital Object Identifier (DOI): see top of this page.
@@ I Introduction
In this work, we tackle the problem of actively reconstructing unknown scenes using posed RGB-D camera data. Given a limited budget, e.g. mission time, our goal is to obtain an accurate 3D representation of the scene by actively positioning the robot with its camera online during a mission. Existing active scene reconstruction frameworks mainly rely on conventional map representations such as voxel grids, meshes, or point clouds [17, 32, 31, 2, 30, 40, 39]. However, these methods often do not deliver high-fidelity reconstruction results due to their sparse representations. Recent advances in implicit neural representations, e.g. neural radiance fields (NeRFs) [21], have attracted significant research interest for their accurate dense scene reconstruction capabilities and low memory footprints. In the context of active reconstruction, several emerging works [25, 5, 13, 14, 27, 37] incorporate uncertainty estimation in NeRFs and exploit it to guide view planning. While these approaches demonstrate promising results, the rather costly volumetric rendering procedure during online incremental mapping poses limitations for NeRF-based active scene reconstruction.
@@
We make the following three claims: (i) our ActiveGS framework achieves superior reconstruction performance compared to state-of-the-art NeRF-based approach and GS-based baselines; (ii) our explicit confidence modelling for Gaussian primitives enables informative viewpoint evaluation and targeted inspection around under-reconstructed surfaces, further improving mission efficiency and reconstruction quality; (iii) we validate our approach in different synthetic indoor scenes and in a real-world scenario using an unmanned aerial vehicle (UAV). To support reproducibility and future research, we open-source our implementation code at: https://github.com/dmar-bonn/active-gs.
@@ II-B Active Scene Reconstruction
To address this, recent research explores implicit neural representations for active reconstruction applications. In an object-centric setup, several methods incorporate uncertainty estimation into NeRFs [25, 13, 14, 37, 27] and use this information to select next best viewpoints. For scene-level reconstruction, Yan et al. [38] and Kuang et al. [18] investigate the loss landscape of implicit neural representations during training to identify under-reconstructed areas. NARUTO [5] learns an uncertainty grid map alongside a hybrid neural scene representation, guiding data acquisition in uncertain regions. These implicit neural representations often face challenges such as inefficient map updates and catastrophic forgetting during incremental mapping.
Several concurrent works propose using GS for active scene reconstruction. GS-Planner [15] and HGS-planner [35] incorporate unknown voxels into the GS rendering pipeline and detect unseen regions for exploration. Li et al. [19] use a Voronoi graph to extract a traversable topological map from the GS representation for path planning. The approach is designed for a 2D planning space, reducing its effectiveness in cluttered environments. FisherRF [12] evaluates the information content of a novel view by measuring the Fisher information value in the GS parameters. This procedure requires computationally expensive gradient calculations at each previously visited and candidate viewpoint, making view planning inefficient for online missions. We build upon the idea of using GS for active scene reconstruction, while introducing a key innovation: we explicitly model the confidence of each Gaussian primitive, enabling viewpoint sampling around low-confidence Gaussian primitives for targeted inspection and fast feed-forward confidence rendering for efficient viewpoint evaluation.
@@ III Our Approach
Figure 2: An overview of our framework. Our hybrid map representation consists of a 2D GS map for high-fidelity scene reconstruction with a coarse voxel map for exploration and path planning. Our view planner leverages unexplored regions in the voxel map for exploration and low-confidence Gaussian primitives for targeted inspection, collecting informative measurements at planned viewpoints for map updates. We iterate between the map update and view planning steps until the pre-allocated mission time is reached.
@@ III-E Viewpoint Sampling and Evaluation
Our viewpoint sampling strategy involves two types of candidate viewpoints. First, we randomly sample $N_{\text{random}}$ candidate viewpoints within a specified range around the current viewpoint. However, relying solely on random local sampling can lead to local minima. To address this, we introduce additional candidate viewpoints based on regions of interest (ROI) defined in the voxel map. We begin by identifying frontier voxels [36] and add them to our ROI set $\mathcal{R}$. By explicitly modelling the confidence of each Gaussian primitive, we can identify and also include voxels containing low-confidence Gaussian primitives in $\mathcal{R}$. Inspired by previous work [17], we define normals for each voxel in $\mathcal{R}$ to indicate the most informative viewing direction. For voxels with low-confidence Gaussian primitives, this is simply the average normal of these Gaussian primitives. The normal of frontier voxels is determined by finding their neighbouring free voxels and calculating the average directional vector from the frontier voxel to these neighbours. To generate ROI-based candidate viewpoints, we create a fixed number of candidate viewpoints within the cone defined by the minimum and maximum sampling distances to each ROI’s centre, and the maximum angular difference relative to its normal. Starting from the closest ROI voxel, we continue outward until we have collected up to $N_{\text{ROI}}$ viewpoints in free space. We illustrate the sampling process in Fig. 3.
@@ IV-A Implementation Details
We test our implementation on a desktop PC with an Intel Core i9-10940X CPU and an NVIDIA RTX A5000 GPU. In this setup, one mapping and planning steps take approximately $1$ s and $0.5$ s, respectively. The whole framework consumes $4-5$ GB GPU RAM during an online mission, with approximately $10$% allocated to the voxel map update.
@@ IV-B Simulation Experiments
  • Ours (w/o ROI): A variant of our ActiveGS that leverages only local random sampling, with $N_{\text{ROI}}=0$.
  • Ours: A variant of our ActiveGS with an alternative confidence formulation, assigning higher confidence to Gaussian primitives with more visible viewpoints, without considering their spatial distribution.
  • FBE [36]: Frontier-based exploration framework that solely focuses on covering unexplored regions, without accounting for the quality of the GS map. We use the collected RGB-D data to update the GS map, similar to our framework.
@@
We present the results of simulation experiments in Fig. 4. Our approach achieves the best performance in both rendering and mesh quality across all test scenes, supporting our first claim that it outperforms state-of-the-art NeRF and GS-based methods. The NeRF-based active scene reconstruction framework, NARUTO, exhibits a significant performance gap compared to our approach, particularly in RGB rendering. This disparity arises because NeRF-based methods often compromise model capacity for faster map updates, limiting their representation quality in scene-level reconstruction. FisherRF evaluates viewpoint utility by calculating the Fisher information in the parameters of the Gaussian primitives within its field of view. This requires computationally expensive gradient calculation for all candidate and training viewpoints, leading to prolonged planning times and incomplete reconstruction under limited mission time. Additionally, since Fisher information is conditioned on the candidate viewpoint, the viewpoint must be selected before its utility can be evaluated, preventing direct viewpoint sampling informed by Fisher information. In contrast, our approach models the confidence of each Gaussian primitive, enabling fast feed-forward confidence rendering for viewpoint evaluation and identification of low-confidence surfaces for targeted candidate viewpoint generation, significantly enhancing reconstruction quality and efficiency. FBE focuses solely on exploration and ignores surface reconstruction quality, limiting its performance. While our approach balances exploration and exploitation by accounting for both unexplored regions and low-confidence Gaussian primitives. The ablation study comparing Ours and Ours (w/o ROI) demonstrates the benefits of ROI-based sampling for targeted inspection, reflected by higher means and smaller standard deviations in both evaluation metrics. Our confidence formulation also outperforms the variant in Ours by considering viewpoint distribution. These results confirm that our confidence modelling is effective in achieving efficient and high-fidelity active scene reconstruction, validating our second claim. We visually compare the reconstruction results in Fig. 5.
@@ IV-C Real-World Experiments
We demonstrate the applicability our framework in a real-world experiment using a UAV equipped with an Intel RealSense 455 RGB-D camera to reconstruct a scene of size $6$ m $\times$ $6$ m $\times$ $3$ m. Unlike simulation experiments, we do not account for the pitch angle of viewpoints in this experiment due to control limitations. The UAV pose is tracked by an OptiTrack motion capture system. Given the limited on-board resources, we run ActiveGS on our desktop PC, where it receives RGB-D and pose data from the UAV for map updates and sends planned collision-free waypoints to guide the UAV. All communication is handled via ROS [26].
@@ V Conclusions and Future Work
A limitation of our current framework is the assumption of perfect localisation. Future work will incorporate localisation uncertainty into confidence modelling of Gaussian primitives. We also plan to better integrate the voxel map with GS map for more efficient mapping and leverage optimisation-based approaches to enhance view planning quality.